[Brahms-l] Shift report 20040119 00:00-08:00

From: Apache <apache@bnl.gov>
Date: Mon Jan 19 2004 - 08:09:11 EST
------------------------------------------------------------
Date: Mon Jan 19 08:09:10 2004, Shift: 00:00-08:00
Supervisor: ZIS, with: 
Log book#15 pages 37 - 37

Shift Summary:
------------------------------------------------------------
1.Beam from the previos shift until 0:47, and the from 02:00 until the end of the shift.

2. Data runs:
   9759 - 161K @trig6, 43K @trig3 (MRS 90, FS 12, FS12)  
   9760 - 346K @trig6, 91K @trig3 (MRS 90, FS 12, FS12)
   9761 - 273K @trig6, 71K @trig3 (MRS 90, FS 12, FS12)
   9762 - 186K @trig6, 36K @trig3 (MRS 90, FS 10, FS10) 
   9763 - 121K @trig6, 23K @trig3 (MRS 90, FS 10, FS10)
3. -FS and BF rotated from  12 to 10.
   -H1 top 16 is out. Need to check if it is a dsic. 
   or a tdc  Adc looks fine.(db) 
   -DVM for t1 and t2 are not working. 
    The Hv is on so not sure what the problem is. (db) 
------------------------------------------------------------

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Mon Jan 19 08:09:22 2004

This archive was generated by hypermail 2.1.8 : Mon Jan 19 2004 - 08:09:40 EST